### v5

# Fusion-PSRO: Nash Policy Fusion for Policy Space Response Oracles

Fusion-PSRO: Nash Policy Fusion for Policy Space Response Oracles
[…]
4998
Appendix A Algorithm for Fusion-PSD-PSRO
Algorithm 1 Fusion-PSD-PSRO
1:Input: initial policy sets for all players $\Pi$
2:Compute utilities $U^{\Pi}$ for each joint $\pi\in\Pi$
3:Initialize Meta-NE $\sigma_{i}=\text{UNIFORM}(\Pi_{i})$
4:for e $\in\{1,2,\dots\}$ do
5:     for player $i\in\{1,2,\dots,n\}$ do
6:         Initialize $\pi^{t+1}_{i}$ via Algorithm LABEL:Fusion and sample $J$ policies $\{\pi^{j}_{i}\}^{J}_{j=1}$ from Policy Hull $\Pi^{t}_{i}$
7:         for many episodes do
8:              Sample $\pi_{-i}\sim\sigma^{t}_{-i}$ and collect the trajectory $\tau$ by playing $\pi_{i}$ against $\pi_{-i}$
9:              Discount the terminal reward $u(\pi_{i},\pi_{-i})$ to each state as the extrinsic reward $r_{1}$
10:              Discount $R^{kl}(\tau)$ to each state as the intrinsic reward $r_{2}$
11:              Store $(s,a,s^{\prime},r)$ to the buffer, where $s^{\prime}$ is the next state and $r=r_{1}+r_{2}$
12:              Estimate the gradient with the samples in the buffer and train oracle $\pi^{t+1}_{i}$ over $\rho\sim(\pi^{t+1}_{i},\pi_{-i})$
13:         end for
14:         $\Pi^{t+1}_{i}=\Pi^{t}_{i}\cup\{\pi^{t+1}_{i}\}$
15:     end for
16:     Compute missing entries in $M^{t+1}$
17:     Compute a Meta-NE $\sigma$ from $M^{t+1}$
18:end for
19:Output: current Meta-NE for each player.
Appendix B Additional Experienments
Figure 1: Average Reward for each iteration (61-100) during the training of each approximate BR within PSRO and PSD-PSRO. Part 1 (top row) shows initial comparisons, while Part 2 (bottom rows) provides detailed comparisons across multiple reward settings.
Appendix C Average Reward during Training on Liar’s Dice
In this ablation experiments, we initialized the initial populations using these four initialization methods respectively and continued using them to generate approximate BRs (corresponding to Normal-BR, Orthogonal-BR, Kaiming-BR and Inherited-BR respectively) during new policy training within PSRO and PSD-PSRO. Simultaneously, we replaced these initialization methods with the Nash Policy Fusion method to generate the second BR (Fusion-BR) for comparison. Fig. 1 shows the average reward for each iteration (61-100) during the training of each approximate BR within PSRO and PSD-PSRO.
Appendix D Benchmark and Implementation Details
D.1 Non-Transitive Mixture Game
This game is characterized by seven Gaussian humps that are evenly spaced on a two-dimensional plane. Each policy in the game corresponds to a point on this plane, analogous to the weights (the probability of that point within each Gaussian distribution) that players allocate to the humps. The payoff containing both non-transitive and transitive components is $\pi^{T}_{i}S\pi_{-i}+\frac{1}{2}\sum^{7}_{k=1}(\pi^{k}_{i}-\pi^{k}_{-i})$, where
$S=\begin{bmatrix}0&1&1&1&-1&-1&-1\\ -1&0&1&1&1&-1&-1\\ -1&-1&0&1&1&1&-1\\ -1&-1&-1&0&1&1&1\\ 1&-1&-1&-1&0&1&1\\ 1&1&-1&-1&-1&0&1\\ 1&1&1&-1&-1&-1&0\\ \end{bmatrix}.$
Contrary to conventional approaches, where PSRO and PSROrN are typically limited by the number of threads of only 1—compared to 4 in other PSRO variants—thus severely constraining their exploration capabilities, we opted for a different way. To mitigate this constraint, we tripled the training episodes for these two algorithms compared to others, aiming to balance the exploration limitations.
Table 1: Hyperparameters for Leduc Poker.
Hyperparameters Value
Oracle
Oracle agent DQN
Replay buffer size $10^{4}$
Mini-batch size 512
Optimizer Adam
Learning rate $5\times 10^{-3}$
Discount factor ($\gamma$) 1
Epsilon-greedy Exploration ($\epsilon$) 0.05
Target network update frequency 5
Policy network MLP (256-256-256)
Activation function in MLP ReLu
PSRO
Episodes for each BR training $2\times 10^{4}$
meta-policy solver Nash
PSD-PSRO
Episodes for each BR training $2\times 10^{4}$
Meta-strategy solver Nash
Diversity weight ($\lambda$) 1
Policy Fusion
Fusion start iteration 2
Fusion method Nash Policy Fusion
Table 2: Hyperparameters for Liar’s Dice.
Hyperparameters Value
Oracle
Oracle agent Rainbow-DQN
Replay buffer size $10^{5}$
Mini-batch size 512
Optimizer Adam
Learning rate $5\times 10^{-4}$
Learning rate decay linear decay
Discount factor ($\gamma$) 0.99
Epsilon-greedy Exploration ($\epsilon$) 0.05
Target network update frequency 5
Network soft update ratio 0.005
Prioritized Experience Replay parameter 0.6
Important sampling parameter 0.4
Gradient clip 10
Policy network MLP (256-256-128)
Activation function in MLP ReLu
PSRO
Episodes for each BR training $2\times 10^{5}$
meta-policy solver Nash
PSD-PSRO
Episodes for each BR training $2\times 10^{5}$
Meta-strategy solver Nash
Diversity weight ($\lambda$) 1
Policy Fusion
Fusion start iteration 2
Fusion method Nash Policy Fusion
Table 3: Hyperparameters for Goofspiel (5 point cards).
Hyperparameters Value
Oracle
Oracle agent DQN
Replay buffer size $10^{4}$
Mini-batch size 512
Optimizer Adam
Learning rate $5\times 10^{-3}$
Discount factor ($\gamma$) 1
Epsilon-greedy Exploration ($\epsilon$) 0.05
Target network update frequency 5
Policy network MLP (512-512-512)
Activation function in MLP ReLu
PSRO
Episodes for each BR training $3\times 10^{4}$
meta-policy solver Nash
PSD-PSRO
Episodes for each BR training $3\times 10^{4}$
meta-strategy solver Nash
diversity weight ($\lambda$) 1
Policy Fusion
Start fusion iteration 20
Fusion method Nash Policy Fusion
D.2 Leduc Poker
Since Diverse-PSRO cannot scale to the RL setting and the code for BD&RD-PSRO in complex games is unavailable, we compare PSRO and PSD-PSRO with their policy fusion counterparts, Fusion-PSRO and Fusion-PSD-PSRO. We implement the PSRO paradigm with Meta-Nash solver, using DQN as the oracle agent. Hyper-parameters are shown in Table 1.
D.3 Liar’s Dice
The Liar’s Dice game involves two players, each equipped with a single die. Incorrect challenges or bids result in immediate loss of the game, as they lead to the loss of dice. We implement the PSRO paradigm with Nash solver, using Rainbow-DQN as the oracle agent. Hyper-parameters are shown in Table 2, and the same set of parameters is used for Liar’s Dice IR as well.
D.4 Goofspiel
In Goofspiel, we use DQN as the oracle agent. Hyper-parameters are shown in Table 3. The start fusion iteration is set to 20.
Figure 2: Exploitability Comparison: Fusion vs. Random Inheritance in Goofspiel (PSD-PSRO)
D.5 Statistical Comparison with Random Inheritance
The following analysis compares exploitability metrics between random inheritance baselines and our fusion methods in two imperfect-information games. Lower exploitability values indicate stronger strategy robustness.
Table 4: Exploitability Statistics for Liar’s Dice (100 iterations)
Method Mean Max Min
Random Inherit (PSRO) 0.0474 0.0556 0.0393
Fusion-PSRO 0.0107 0.0175 0.0039
Random Inherit (PSD) 0.0376 0.0411 0.0341
Fusion-PSD-PSRO 0.0148 0.0264 0.0045
Table 4 shows consistent superiority of fusion methods in Liar’s Dice. Both Fusion-PSRO and Fusion-PSD-PSRO achieve substantially lower exploitability across all metrics compared to their random inheritance counterparts, demonstrating enhanced strategy robustness.
Table 5: Exploitability Statistics for Goofspiel (160 iterations)
Method Mean Max Min
Random Inherit (PSD-PSRO) 0.4836 0.5332 0.4340
Fusion-PSD-PSRO 0.4004 0.4127 0.3881
As evidenced in Table 5, the fusion approach maintains its advantage in Goofspiel, with notably reduced exploitability values and tighter performance ranges compared to random inheritance. This pattern is further illustrated in Figure 2, showing the evolutionary trajectories of both methods.
D.6 Time Consumption in Fusion-PSRO
We analyze the time consumption specific to Nash Policy Fusion within the Fusion-PSRO framework. As shown in Table 6, the computation of the fusion policy parameters accounts for only 0.01% of the total time during training on Leduc Poker. This indicates that Nash Policy Fusion is an efficient addition to the PSRO without significantly impacting overall runtime.
Main Part of Fusion-PSRO Percentage (%)
Compute NE of the meta game 0.19
Compute the Approximate BR 81.85
Compute the Fusion Policy Parameters 0.01
Compute the Payoff Matrix 17.69
Table 6: Time Consumption in Fusion-PSRO on Leduc Poker.